OpenStudy (anonymous):

how to find the Slope of -2x+y=4??

OpenStudy (anonymous):

Change this into y=Mx+C form. Where m is the slope and c is constant. so slope is 2

OpenStudy (anonymous):

-2x + y = 4 or, y = 2x + 4 So, the slope is 2. Remember, for any straight line of the form, y = mx + c m is the slope and c is the intercept.

OpenStudy (anonymous):

Y=mx+b You mean??

OpenStudy (anonymous):

y = mx + c this is the general equation for a line.

OpenStudy (anonymous):

how do you graph this?

OpenStudy (anonymous):

you can graph the line by putting the constant values in X and Y say example if x=0 then point y=4..that is how it works